A JAZZ CONCERT BROUGHT IN $164,000 ON THE SALE OF 7,800 TICKETS. IF THE TICKETS SOLD FOR $15 AND $25 EACH, HOW MAY OF EACH TYPE OF TICKET WERE SOLD?
Answer by mananth(16949)

x= Number of $25 tickets
y= Number of $15 tickets

1 x + 1 y = 7800 .............1
Total value
16 x + 25 y = 164000 .............2
Eliminate y
multiply (1)by -25
Multiply (2) by 1
-25 x -25 y = -195000
16 x + 25 y = 164000
Add the two equations
-9 x = -31000
/ -9
x = 3444
plug value of x in (1)
1 x + 1 y = 7800
3444 + y = 7800
y = 7800 -3444
y = 4356
y = 4356
x= 3444 Number of $25 tickets
y= 4356 Number of $15 tickets
m.ananth@hotmail.ca